On planet X, an object falls 18 feet in 3 seconds. Knowing the distance it falls varies directly with the square of the time of fall, how long does it take an object to fall 100 feet? Round to three decimal places, if necessary.
Answer by stanbon(48558)

On planet X, an object falls 18 feet in 3 seconds. Knowing the distance it falls varies directly with the square of the time of fall, how long does it take an object to fall 100 feet? Round to three decimal places, if necessary.
d=kt^2; need to find "k"
18=k(3^2)
k=6
Therefore d=6t^2 on that planet.
If d=100 then 100=6t^2
t^2=100/6
t=[10sqrt6]/6 = 4.082... seconds
Cheers,
Stan H.